# Optimal Estimation of Temperature

### v4

@@ Appendix C Relative biases of $\beta$ and $T$ in a two-level system
$\begin{split}\Phi_{T}&=\left|\frac{\langle\hat{\beta}^{(-1)}\rangle-T}{T}\right|\\ &=\frac{2^{N}\text{e}^{\frac{-\beta N\varepsilon}{2}}}{(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}})^{N}}\\ &=\left(\frac{2\text{e}^{\frac{-\beta\varepsilon}{2}}}{\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}}\right)^{N}\\ &=\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}}{\cosh(\frac{\beta\varepsilon}{2})^{N}}\\ &=\text{e}^{-N(\frac{\beta\varepsilon}{2}+\ln\cosh(\frac{\beta\varepsilon}{2}))}\\ &\sim\text{e}^{-\alpha_{2}N}.\end{split}$ (41)
In FIG. 6b, when $T<0$, $\alpha_{2}$ becomes negative, causing $\Phi_{T}$ to increase exponentially, and the unbiased estimator of $\hat{\beta}^{(-1)}$ becomes invalid. When $T>0$, $\alpha_{2}$ is positive, and the unbiased estimator of $\hat{\beta}^{(-1)}$ is valid, except at extremely high temperatures where $\alpha_{2}$ approaches zero.
@@ Appendix D Unbiased estimator of $\beta^{\alpha}$ for $\alpha\in\mathbb{R}$
In this appendix, let us consider the unbiased estimation of the non-integer powers of temperature, i.e., $\langle\hat{\beta}^{(\alpha)}\rangle=\beta^{\alpha}$ for $\alpha\in R$. The expression of $\hat{\beta}^{(\alpha)}$ is still given by Eq. (6) except that we extend the definition of $\text{d}^{\alpha}_{E}\sigma(E)$ in Eq. (7) in different cases. When $\alpha<0$ , the definition is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{E}\sigma(u)\left(E-t\right)^{-\alpha-1}\text{d}u,$ (42)
@@
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha-1}\\ &=\frac{1}{Z}\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{-\alpha-1}\\ &=\frac{1}{Z}\frac{\beta^{\alpha}}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha-1}\\ &=\frac{1}{Z}\beta^{\alpha}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\\ &=\beta^{\alpha}.\end{split}$ (43)
When $\alpha>0$ and $n-1<\alpha<n$ for some integer $n$, the definition is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{E}\left(E-t\right)^{n-\alpha-1}\text{d}^{n}_{u}\sigma(u)\text{d}u,$ (44)
@@
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle=&\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ =&\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ =&\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(n-\alpha)}\\ &\int_{E_{\text{min}}}^{E}\text{d}t\text{d}_{t}^{n}\sigma(t)\left(E-t\right)^{n-\alpha-1}\\ =&\frac{1}{Z}\frac{1}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\\ &\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{n-\alpha-1}\\ =&\frac{1}{Z}\frac{\beta^{\alpha-n}}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha-1}\\ =&\frac{1}{Z}\beta^{\alpha-n}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\text{e}^{-\beta t}\\ =&\beta^{\alpha}.\end{split}$ (45)
The last equlity is obtained by using Eq. (29).
When $\alpha>0$ and the condition $\left.\text{d}_{E}^{l}\sigma(E)\right|_{E=0}=0$ is not satisfied for every $l\in\left[0,n-1\right]$, we only consider the case $\alpha\in\left[0,1\right)$. Here, the definition is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(1-\alpha)}\frac{\text{d}}{\text{d}E}\int_{E_{\text{min}}}^{E}\left(E-t\right)^{-\alpha}\sigma(u)\text{d}u,$ (46)
@@
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle&=\frac{1}{Z}\sum_{i\in I}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(1-\alpha)}\\ &\quad\frac{\text{d}}{\text{d}E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\left.\frac{1}{\Gamma(1-\alpha)}\text{e}^{-\beta E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\right|^{\infty}_{0}\\ &\quad-\frac{1}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}\text{e}^{-\beta E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\frac{\beta}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\frac{\beta^{\alpha}}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha}\\ &=\frac{1}{Z}\beta^{\alpha}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\\ &=\beta^{\alpha}.\end{split}$ (47)
@@ Appendix E The approximate expression of $\hat{\beta}_{B}$ and $\hat{T}_{G}$ in large-$N$ limit
Using the Laplace approximation [72, 73] or the Darwin–Fowler method in Sec. 9.1 of [67], $\ln Z$ can be approximated as [84]
$\begin{split}\ln Z=&\ln\sigma(E^{*})+\ln\sqrt{\left.2\pi\left(-\frac{\partial\hat{\beta}_{B}}{\partial E}\right)^{-1}\right|_{E=E^{*}}}\\ &-\beta E^{*}+O\left(N^{-1}\right),\end{split}$ (48)
@@
$\left\langle E\right\rangle=E^{*}+\left.\frac{1}{2}\frac{\partial^{2}\hat{\beta}_{B}}{\partial E^{2}}\right|_{E=E^{*}}\left(\frac{\partial E^{*}}{\partial\beta}\right)^{2}+O(N^{-1}).$ (49)
@@
$E^{*}=\left\langle E\right\rangle-T+\frac{T^{2}}{2}\frac{\partial\ln C}{\partial T}+O(N^{-1}),$ (50)
where $C=\partial\left\langle E\right\rangle/\partial T$ denotes the heat capacity. Substituting Eq. (50) into Eq. (48), we have
where $S=\ln Z-\beta\partial\ln Z/\partial\beta$ is the thermodynamics entropy in a canonical ensemble. Eq. (51) shows the nonequivalence between micro-canonical ensemble and canonical ensemble for finite-$N$ systems, i.e., the relation between the Boltzmann entropy $S_{B}=\ln[\sigma(E)\epsilon]$ and the thermodynamic entropy $S(E)$. If we replace $\left\langle E\right\rangle$ with a sample of system’s energy $E$, the function $\hat{\beta}_{L}(E)$ (given by $\left.\frac{\partial\ln Z}{\partial\beta}\right|_{\beta=\hat{\beta}_{L}(E)}=E$) will be a statistic of $\beta$ [85]. Taking the derivative of Eq. (51) with respect to $E$ on both sides, we have
$\hat{\beta}_{B}(E)=\hat{\beta}_{L}(E)+\frac{1}{2}\frac{\text{d}^{2}_{E}\hat{\beta}_{L}(E)}{\text{d}_{E}\hat{\beta}_{L}(E)}+O(N^{-2}).$ (52)
@@
Performing Taylor expansion of Eq. (52) at $\langle E\rangle$ with $\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)=\beta$, and considering $\delta E=E-\langle E\rangle$ as a small quantity, we have
@@
$\begin{split}\hat{\beta}_{B}-\beta=&\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\delta E+\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\frac{(\delta E)^{2}}{2}\\ &+\text{d}_{\left\langle E\right\rangle}^{3}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\frac{(\delta E)^{3}}{6}+\frac{1}{2}\frac{\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}\\ &+\text{d}_{\left\langle E\right\rangle}\left(\frac{\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}\right)\frac{\delta E}{2}+O(N^{-2}).\end{split}$ (53)
@@
Using Eq. (53), we obtain the variance of $\hat{\beta}_{B}$, $\Delta\beta^{2}_{B}=\langle(\hat{\beta}_{B}-\beta)^{2}\rangle$, which can be expanded as
@@
$\begin{split}\Delta\beta^{2}_{B}=&\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)^{2}\left\langle(\delta E)^{2}\right\rangle+\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left(\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left\langle(\delta E)^{3}\right\rangle\\ &+\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\text{d}_{\left\langle E\right\rangle}\left(\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\right)\left\langle(\delta E)^{2}\right\rangle\\ &+\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{2}\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\left\langle(\delta E)^{2}\right\rangle+\frac{1}{4}\left(\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\right)^{2}\\ &+\frac{\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left(\text{d}^{3}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)}{3}\left\langle(\delta E)^{4}\right\rangle+\frac{\left(\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)^{2}}{4}\left\langle(\delta E)^{4}\right\rangle\\ &+O(N^{-3}).\end{split}$ (54)
@@
In the derivation, we have used $\delta E\sim N^{1/2}$ and $\text{d}_{\left\langle E\right\rangle}\sim N^{-1}$.
To simplify Eq. (54), we introduce the higher order derivatives of the inverse function [86],

### v5

@@ C.2 Relative bias of $T$ in a two-level system
$\begin{split}\Phi_{T}&=\left|\frac{\langle\hat{\beta}^{(-1)}\rangle-T}{T}\right|\\ &=\frac{2^{N}\text{e}^{\frac{-\beta N\varepsilon}{2}}}{(\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}})^{N}}\\ &=\left(\frac{2\text{e}^{\frac{-\beta\varepsilon}{2}}}{\text{e}^{\frac{\beta\varepsilon}{2}}+\text{e}^{\frac{-\beta\varepsilon}{2}}}\right)^{N}\\ &=\frac{\text{e}^{\frac{-\beta N\varepsilon}{2}}}{\cosh(\frac{\beta\varepsilon}{2})^{N}}\\ &=\text{e}^{-N(\frac{\beta\varepsilon}{2}+\ln\cosh(\frac{\beta\varepsilon}{2}))}\\ &\sim\text{e}^{-\alpha_{2}N}.\end{split}$ (C7)
In figure 6b, when $T<0$, $\alpha_{2}$ becomes negative, causing $\Phi_{T}$ to increase exponentially, and the unbiased estimator of $\hat{\beta}^{(-1)}$ becomes invalid. When $T>0$, $\alpha_{2}$ is positive, and the unbiased estimator of $\hat{\beta}^{(-1)}$ is valid, except at extremely high temperatures where $\alpha_{2}$ approaches zero.
@@ Appendix D Unbiased estimator of $\beta^{\alpha}$ for $\alpha\in\mathbb{R}$
In this appendix, let us consider the unbiased estimation of the non-integer powers of temperature, i.e., $\langle\hat{\beta}^{(\alpha)}\rangle=\beta^{\alpha}$ for $\alpha\in R$. The expression of $\hat{\beta}^{(\alpha)}$ is still given by Eq. (6) except that we extend the definition of $\text{d}^{\alpha}_{E}\sigma(E)$ in Eq. (7) in different cases.
D.1 Unbiased estimator of $\beta^{\alpha}$ for $\alpha<0$
When $\alpha<0$ , it is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{E}\sigma(u)\left(E-t\right)^{-\alpha-1}\text{d}u,$ (D1)
@@ D.1 Unbiased estimator of $\beta^{\alpha}$ for $\alpha<0$
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle&=\frac{1}{Z}\sum_{i\in D}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha-1}\\ &=\frac{1}{Z}\frac{1}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{-\alpha-1}\\ &=\frac{1}{Z}\frac{\beta^{\alpha}}{\Gamma(-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha-1}\\ &=\frac{1}{Z}\beta^{\alpha}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\\ &=\beta^{\alpha}.\end{split}$ (D2)
D.2 Unbiased estimator of $\beta^{\alpha}$ for $\alpha>0$ and $n-1<\alpha<n$
When $\alpha>0$ and $n-1<\alpha<n$ for some integer $n$, it is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{E}\left(E-t\right)^{n-\alpha-1}\text{d}^{n}_{u}\sigma(u)\text{d}u,$ (D3)
@@ D.2 Unbiased estimator of $\beta^{\alpha}$ for $\alpha>0$ and $n-1<\alpha<n$
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle=&\frac{1}{Z}\sum_{i\in D}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ =&\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ =&\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{E}\text{d}t\text{d}_{t}^{n}\sigma(t)\left(E-t\right)^{n-\alpha-1}\\ =&\frac{1}{Z}\frac{1}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{n-\alpha-1}\\ =&\frac{1}{Z}\frac{\beta^{\alpha-n}}{\Gamma(n-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha-1}\\ =&\frac{1}{Z}\beta^{\alpha-n}\int_{E_{\text{min}}}^{\infty}\text{d}t\text{d}_{t}^{n}\sigma(t)\text{e}^{-\beta t}\\ =&\beta^{\alpha}.\end{split}$ (D4)
The last equlity is obtained by using Eq. (A2).
D.3 Unbiased estimator of $\beta^{\alpha}$ for $\alpha\in\left[0,1\right)$
When $\alpha>0$ and the condition $\left.\text{d}_{E}^{l}\sigma(E)\right|_{E=0}=0$ is not satisfied for every $l\in\left[0,n-1\right]$, we only consider the case $\alpha\in\left[0,1\right)$. Here, it is extended to
$\text{d}^{\alpha}_{E}\sigma(E)=\frac{1}{\Gamma(1-\alpha)}\frac{\text{d}}{\text{d}E}\int_{E_{\text{min}}}^{E}\left(E-t\right)^{-\alpha}\sigma(u)\text{d}u,$ (D5)
@@ D.3 Unbiased estimator of $\beta^{\alpha}$ for $\alpha\in\left[0,1\right)$
$\begin{split}\langle\hat{\beta}^{(\alpha)}\rangle&=\frac{1}{Z}\sum_{i\in D}\frac{1}{\sigma(E_{i})}\text{d}^{\alpha}_{E_{i}}\sigma(E_{i})d_{i}\text{e}^{-\beta E_{i}}\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\text{d}_{E}^{\alpha}\sigma(E)\\ &=\frac{1}{Z}\int_{E_{\text{min}}}^{\infty}\text{d}E\text{e}^{-\beta E}\frac{1}{\Gamma(1-\alpha)}\frac{\text{d}}{\text{d}E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\left.\frac{1}{\Gamma(1-\alpha)}\text{e}^{-\beta E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\right|^{\infty}_{0}-\frac{1}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}\text{e}^{-\beta E}\int_{E_{\text{min}}}^{E}\text{d}t\sigma(t)\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\frac{\beta}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\int_{t}^{\infty}\text{d}E\text{e}^{-\beta E}\left(E-t\right)^{-\alpha}\\ &=\frac{1}{Z}\frac{\beta^{\alpha}}{\Gamma(1-\alpha)}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\int_{E_{\text{min}}}^{\infty}\text{d}x\text{e}^{-x}x^{-\alpha}\\ &=\frac{1}{Z}\beta^{\alpha}\int_{E_{\text{min}}}^{\infty}\text{d}t\sigma(t)\text{e}^{-\beta t}\\ &=\beta^{\alpha}.\end{split}$ (D6)
@@ Appendix E The approximate expression of $\hat{\beta}_{B}$ and $\hat{T}_{G}$ in large-$N$ limit
Using the Laplace approximation [63, 64] or the Darwin–Fowler method in Sec. 9.1 of [57], $\ln Z$ can be approximated as  [^5] We assume that the system does not undergo a phase transition as the requirement of the Laplace approximation.
$\begin{split}\ln Z=&\ln\sigma(E^{*})+\ln\sqrt{\left.2\pi\left(-\frac{\partial\hat{\beta}_{B}}{\partial E}\right)^{-1}\right|_{E=E^{*}}}-\beta E^{*}+O\left(N^{-1}\right),\end{split}$ (E1)
@@
$\left\langle E\right\rangle=E^{*}+\left.\frac{1}{2}\frac{\partial^{2}\hat{\beta}_{B}}{\partial E^{2}}\right|_{E=E^{*}}\left(\frac{\partial E^{*}}{\partial\beta}\right)^{2}+O(N^{-1}).$ (E2)
@@
$E^{*}=\left\langle E\right\rangle-T+\frac{T^{2}}{2}\frac{\partial\ln C}{\partial T}+O(N^{-1}),$ (E3)
where $C=\partial\left\langle E\right\rangle/\partial T$ denotes the heat capacity. Substituting Eq. (E3) into Eq. (E1), we have
$\ln\sigma(\left\langle E\right\rangle)=S-\ln\sqrt{2\pi\left(-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)}+O(N^{-1}),$ (E4)
where $S=\ln Z-\beta\partial\ln Z/\partial\beta$ is the thermodynamics entropy in a canonical ensemble. Eq. (E4) shows the nonequivalence between micro-canonical ensemble and canonical ensemble for finite-$N$ systems, i.e., the relation between the Boltzmann entropy $S_{B}=\ln[\sigma(E)\epsilon]$ and the thermodynamic entropy $S(E)$. If we replace $\left\langle E\right\rangle$ with a sample of system’s energy $E$, the function $\hat{\beta}_{L}(E)$ (given by $\left.\frac{\partial\ln Z}{\partial\beta}\right|_{\beta=\hat{\beta}_{L}(E)}=E$) will be a statistic of $\beta$  [^6] Note that Eq. (E4) is invalid for $E=0$, since the right-hand side of Eq. (E4) diverges while the left-hand side of Eq. (E4) is a constant. The breakdown of Eq. (E4) is consistent with the fact that $\Delta E^{2}=CT^{2}\to 0$ as $E\to 0$.. Taking the derivative of Eq. (E4) with respect to $E$ on both sides, we have
$\hat{\beta}_{B}(E)=\hat{\beta}_{L}(E)+\frac{1}{2}\frac{\text{d}^{2}_{E}\hat{\beta}_{L}(E)}{\text{d}_{E}\hat{\beta}_{L}(E)}+O(N^{-2}).$ (E5)
@@
Performing Taylor expansion of Eq. (E5) at $\langle E\rangle$ with $\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)=\beta$, and considering $\delta E=E-\langle E\rangle$ as a small quantity, we have
@@
$\begin{split}\hat{\beta}_{B}-\beta=&\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\delta E+\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\frac{(\delta E)^{2}}{2}+\text{d}_{\left\langle E\right\rangle}^{3}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)\frac{(\delta E)^{3}}{6}\\ &+\frac{1}{2}\frac{\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}+\text{d}_{\left\langle E\right\rangle}\left(\frac{\text{d}_{\left\langle E\right\rangle}^{2}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\left(\left\langle E\right\rangle\right)}\right)\frac{\delta E}{2}+O(N^{-2}).\end{split}$ (E6)
@@
Using Eq. (E6), we obtain the variance of $\hat{\beta}_{B}$, $\Delta\beta^{2}_{B}=\langle(\hat{\beta}_{B}-\beta)^{2}\rangle$, which can be expanded as
@@
$\begin{split}\Delta\beta^{2}_{B}=&\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)^{2}\left\langle(\delta E)^{2}\right\rangle+\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left(\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left\langle(\delta E)^{3}\right\rangle\\ &+\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\text{d}_{\left\langle E\right\rangle}\left(\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\right)\left\langle(\delta E)^{2}\right\rangle+\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{2}\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\left\langle(\delta E)^{2}\right\rangle\\ &+\frac{1}{4}\left(\frac{\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}}{\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}}\right)^{2}+\frac{\left(\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)\left(\text{d}^{3}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)}{3}\left\langle(\delta E)^{4}\right\rangle+\frac{\left(\text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}\right)^{2}}{4}\left\langle(\delta E)^{4}\right\rangle+O(N^{-3}).\end{split}$ (E7)
@@
In the derivation, we have used $\delta E\sim N^{1/2}$ and $\text{d}_{\left\langle E\right\rangle}\sim N^{-1}$. To simplify Eq. (E7), we introduce the higher order derivatives of the inverse function [74],